OpenStudy (anonymous):

Apples cost $12 per dozen more than oranges and $18 per dozen less than pears. The cost of 2 oranges and 4 apples is the same as the cost of 4 pears. What is the cost of a pear? @hartnn

OpenStudy (anonymous):

@Directrix @terenzreignz

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what equations have you tried to develop?

OpenStudy (anonymous):

i let $x be the cost of one apple. 2(x-12/12)+4x=4(x+18/12)

OpenStudy (rina.r):

the question is not clear.

OpenStudy (anonymous):

i do think so

OpenStudy (amistre64):

Apples cost: $12 per dozen more than oranges and $18 per dozen less than pears. 12a = 12o+12, solve for o 12a = 12p+16, solve for p 2o +4a = 4p What is p?

OpenStudy (amistre64):

fingers aint awake yet ... 2nd equation: 12a = 12p - 18

OpenStudy (anonymous):

why so complicated @@

OpenStudy (amistre64):

thats not complicated, thats just my interpretation of the information

OpenStudy (anonymous):

my brain hurts if my answer is correct, these pears are way above my price range

OpenStudy (anonymous):

what's your answer?

OpenStudy (amistre64):

a = o+1 2a = 2p-3 o +2a = 2p p = 1/2 o +a p = 1/2 o +o+1 p = 3/2o + 1 o = p - 5/2 p = 3/2 p - 15/4 + 1 -1/2p = -2.75 p = 5.50, maybe?

OpenStudy (anonymous):

that's correct. i think of this question about an hour. omg

OpenStudy (anonymous):

but the topic is "linear equation in one unknown"

OpenStudy (amistre64):

then there is most likely a way to express it as 1 unknown; i got it down to p alone

OpenStudy (anonymous):

ohh.... o=(12a-12)/12 may be use the formula in this way...

OpenStudy (amistre64):

maybe .... need more MtDew 8/

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

o=(12a-12)/12 p=(12a-16)/12 then 2[(12a-12)/12]+4a=4[(12a-16)/12] is it?

OpenStudy (phi):

It does not make sense that you treat this as "linear equation in one unknown" it is 3 variables and 3 equations... and once you get to 3 or more variables I think elimination is the more straight-forward approach. you would solve (in the order a, o, p) 1 -1 0 1 1 0 -1 -1.5 4 2 -4 0

OpenStudy (anonymous):

but this worksheet is focus on linear equation in one unknown....

OpenStudy (anonymous):

why i got -ve number ????

OpenStudy (amistre64):

12a = 12o+12 12a = 12p-18; a = (2p-3)/2 12o + 12 = 12p - 18 2o+2 = 2p-3; o = (2p-5)/2 2a + 4o = 4p; sub in a and o to get it all in terms of p 2(2p-3)/2 +4(2p-5)/2 = 4p 2p-3 +4p-10 = 4p ; solve for p 6p-13 = 4p 2p = 13 p = 13/2 = 6.5 .... of course i most likely got some error in that lol

OpenStudy (anonymous):